When body of mass m is suspended from a spiral spring of natural length L the spring gets stretched through a distance h. The potential energy of stretched spring is?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Spring force is given as:
F = -kx
Work done = F x displacement => Fdx
dW = -kxdx
integerating both sides,
W = -kx^2/2 {integerating from (h  + L) to L.
W = -k/2[(h+L)^2 – L^2]
    = -k/2[h^2 *(h+2L)] Joule.